Sora Industries currently has 60 million outstanding shares, $120 million in debt and $40 million in cash. Sales in the last financial year were $433. They are projected to grow at 8.1% next year, 10.3% in the second year, and 6% in the third year. From the fourth year onwards, sales will grow at 5% per year. Furthermore, you are given that in the last financial year, Sora’s i) cost of goods sold was 67% of sales, ii) selling, general and administrative expenses were 20% of sales, iii) depreciation was 1.5% of sales, iv) gross fixed assets were 22.5% of sales, and v) NWC was 18% of sales. Assuming these ratios remain constant forever, the tax rate for the firm is 40%, and the WACC is 10%:

  1. What is the stock price of Sora Industries?
  2. Sora’s cost of goods sold was assumed to be 67% of sales. If its cost of goods sold is actually 70% of sales, how would the estimate of the stock’s value change?
  3. Sora’s net working capital needs were estimated to be 18% of sales (their current level in year 0). If Sora can reduce this requirement to 12% of sales starting in year 1, but all other assumptions remain as in part (a), what stock price do you estimate for Sora?
Answer #1

year 2 = 46 Current sales = 433 - year i= 433 +8 -1% = 468 Crounded off 2 468 + 10:34 - S16 - 1 3 = 516+ 6%. = 547 to 42 547154.4 103 180.5 189.5 a tarnings and Free cash flow fore cast in $ million I the year No 1 2 3 4 1. sales or 433 468 516 541a WACC= 10% (6) = 252 +246 + 30.8+666 10 Coszt – - 567 (10)} e = 567+40-120 60 4 (6) In case of cost of goods sold @ 10% 7 evТа ині дън Еоіtу bумча 1(3): 22.92 = 458 10% -5%. Vo = 16.88 + 15.32 + 20.94+458 o T (1.1727 (1.13 I = 388 L . Salz - to D.- Every thing remains in Net working capital If we adjust Question (a) to only constant except changes hence It is enough Fre